### v1

# FAM-HRI: Foundation-Model Assisted Multi-Modal Human-Robot Interaction Combining Gaze and Speech

Abstract
Effective Human-Robot Interaction (HRI) is crucial for enhancing accessibility and usability in real-world robotics applications. However, existing solutions often rely on gestures or language commands, making interaction inefficient and ambiguous, particularly for users with physical impairments. In this paper, we introduce FAM-HRI, an efficient multi-modal framework for human-robot interaction that integrates language and gaze inputs via foundation models. By leveraging lightweight Meta ARIA glasses, our system captures real-time multi-modal signals and utilizes large language models (LLMs) to fuse user intention with scene context, enabling intuitive and precise robot manipulation. Our method accurately determines gaze fixation time interval, reducing noise caused by the gaze dynamic nature. Experimental evaluations demonstrate that FAM-HRI achieves a high success rate in task execution while maintaining a low interaction time, providing a practical solution for individuals with limited physical mobility or motor impairments.
{strip}
[…]
figureVoice-gaze fusion HRI enables efficient, intuitive interaction without memorization, ideal for physically impaired users.
Supplementary Material
The supplementary materials and video can be found at: https://github.com/laiyuzhi/FAM-HRI
I INTRODUCTION
Individuals with limited mobility or motor impairments, such as neuromuscular disorders, often struggle with daily tasks. These tasks can be automated using robotic manipulators, enhancing independence and improving quality of life [1, 2, 3]. However, improving life through automation requires effective and accessible Human-Robot Interaction (HRI), which remains a challenge. Ensuring smooth interaction is particularly difficult in systems that rely on verbal commands [4]. Accurately processing verbal descriptions becomes problematic when identifying unfamiliar objects or distinguishing between similar items. Existing voice-controlled assistants, such as Amazon Alexa or Tmall Genie, remain insufficient in bridging this gap for embodied HRI.
Gaze, as a natural interaction modality, has great potential in HRI for individuals with severe physical limitations [5, 6, 7, 8], as shown in Fig. FAM-HRI: Foundation-Model Assisted Multi-Modal Human-Robot Interaction Combining Gaze and Speech. However, existing methods often rely on external or bulky devices, restricting mobility and reducing the accuracy of HRI [5, 7, 8]. Some early gaze-based interfaces [6, 9] enable robot control at the cost of increased interaction time and complexity [10]. Beyond hardware constraints, the nature of gaze detection itself adds challenges [11]. During natural gaze interactions, the gaze does not always remain fixed on the target object but instead briefly lingers around the area [12]. This dynamic nature increases noise in target selection process, reducing gaze stability in daily use. These challenges emphasize the need for accurate, efficient, and lightweight systems to enhance assistive robotics.
To address the challenges, we introduce a Foundation-model Assisted Multi-modal HRI framework, FAM-HRI, that enables users to convey their intentions to robots. By wearing lightweight glasses from the Meta ARIA Research Kit [13], the system can simultaneously collect gaze and voice signals from users. By integrating these modalities, the robot accurately selects target objects and executes intended actions based on environmental context, reducing ambiguity in language-based systems, especially for users with limited mobility. A command-processing LLM analyzes user input to identify the optimal gaze period, extracting high-quality signals while minimizing noise from gaze dynamics. A motion-planning LLM then generates the robot’s action sequences, leveraging advanced reasoning to handle complex contexts more effectively than traditional rule-based systems. Our main contributions are summarized below:
  • To the best of our knowledge, FAM-HRI is the one of the first frameworks to integrate voice and gaze with a lightweight wearable device for human-robot collaboration. By leveraging ARIA glasses, our approach facilitates natural and intuitive communication, enhancing accessibility and ease of use. We assess its effectiveness through a comprehensive user experience study.
  • Our system optimizes gaze fixation timing and robot planning by integrating language, gaze, and environmental context. This ensures precise user intent alignment with robot actions, enhancing task execution.
  • FAM-HRI was benchmarked against state-of-the-art HRI methods, achieving higher interaction success rates and faster input processing. To support the community, we will release our system design, algorithms, and solutions as open-source at github.com/laiyuzhi/FAM-HRI.
II Related Work
With advancements in computer vision and highly integrated hardware, robots are increasingly embedded in daily life, assisting with tasks from home services [14, 15] to industrial automation [16, 17]. To facilitate seamless human-robot collaboration, various interaction modalities have gained attention. However, gesture-based systems [4] and those combining voice and posture [18, 19] impose physical constraints, making them unsuitable for users with limited mobility or motor impairments. In such cases, language-based interaction emerges as a natural alternative. However, language-based control presents challenges, including ambiguity in scene descriptions and multiple interpretations of the same instruction, reducing reliability in complex environments [19]. To address this, some researchers [20] proposed a vision-language model (VLM)-based dialog system, enabling interactive refinement between the robot and user. While this improves robustness, it significantly increases interaction time, limiting real-time efficiency. Alternatively, other studies [21, 22] explored LLMs for robotic task planning using language input. However, LLMs and VLMs lack pretraining in complex geometric reasoning, leading to failures in scenarios with multiple similar objects.
Figure 1: FAM-HRI Framework. Gaze and speech inputs from ARIA glasses are processed by GPT agents to determine the target object and its fixation interval. The human view is aligned with the robot view via feature matching, and a planning policy generates parameterized robot actions.
Among alternative modalities, gaze offers a powerful interaction method for individuals with disabilities. Most existing gaze-based assistive devices rely on graphical user interface (GUI) systems, which enable users to control assistive devices through visual interfaces. Some researchers employ virtual buttons, allowing users to operate electric wheelchairs [9], robotic arms [23], or mobile robots [24]. However, GUI-based systems divert user attention from the external environment, limiting their field of view. An optimal system should enable unrestricted gaze-based interaction by accurately estimating user attention [10]. In [25], researchers introduced a gaze-following model that predicts gaze direction and visual attention using RGB images from a robot’s perspective combined with YOLO-based object detection. Similarly, [26] applies transfer learning to track gaze within predefined regions, allowing the system to determine attention states. This enhances the ability of the robot to interpret user intentions, improving efficiency and naturalness in human-robot collaboration, particularly in industrial settings. However, these approaches also require users to fixate on targets continuously, which feels unnatural compared to typical communication, where the gaze naturally shifts.
Existing HRI approaches remain inadequate for general use, particularly for individuals with difficulties in language commands or physical movement. Integrating LLMs with strong reasoning capabilities offers potential for enhancing multi-modal data processing across various HRI applications, including social robots [18, 19] and robot navigation [27]. Recently, researchers introduced LaMI [28] to explore how LLMs can improve multi-modal HRI by reasoning over diverse inputs. However, this approach primarily focuses on backend reasoning and overlooks frontend measurement uncertainties and other external factors affecting perception and decision-making in real-world interactions.
Unlike previous HRI approaches that struggle with frontend uncertainties and real-world perception challenges, our method leverages lightweight ARIA glasses to capture high-precision gaze data and speech signals. Using LLM-based reasoning, we accurately determine the most probable gaze fixation period, minimizing noise from gaze dynamics. Additionally, a second LLM agent analyzes scene context and user intent to generate optimal robot actions, ensuring more efficient and natural HRI.
III Problem Definition
Our proposed system integrates environmental perception from both the human and the robot points of view to facilitate effective collaboration. The human view is mainly used for intention fusion, capturing user-centric cues, while the robot view focuses on manipulation, allowing precise interaction with objects. To structure multi-modal signals for foundation model reasoning, we need to define sets of geometric frames and control parameters.
III-A Definitions of Frames and Transformations
To ensure consistency, we define the reference frames used in this paper. The robot base frame ()$\prescript{r}{}{(\cdot)}$ serves as the primary reference for the robot, while the robot camera frame ()$\prescript{c}{}{(\cdot)}$ is used for manipulation. The ARIA glasses camera frame (()$\prescript{gc}{}{(\cdot)}$) provides a first-person user-centric perspective. The ARIA pupil frame (()$\prescript{gp}{}{(\cdot)}$), centered at the midpoint of the glasses’ bridge, serves as the reference for gaze vector estimation. The transformation $T\in\mathbb{SE}(3)$ denotes a rigid body transformation in the special Euclidean group. Specifically, the transformation matrix ${}^{gc}T_{gp}$ maps points from the ARIA pupil frame to the ARIA camera frame, with parameters provided by the manufacturer111facebookresearch.github.io/projectaria_tools/.
III-B Formulation of Human View Representation
To formally characterize the human view inputs of ARIA glasses, we define the human input state vector as $\mathcal{Z}_{H}=(\mathcal{S},\mathcal{U},\mathcal{G})$. Here, $\mathcal{S}\in\Sigma^{*}$ denotes the transcribed speech sequence. The audio is recorded at a 48 kHz sampling rate using the built-in microphone of the ARIA glasses and converted into text via Fast Whisper [29], with each word timestamped. Six embedded microphones improve speech clarity by suppressing background noise and improving transcription accuracy even in noisy environments. Additional details are provided in the supplementary material. $\mathcal{U}\in\mathbb{R}^{H_{1}\times W_{1}\times(t_{n}-t_{1})}$ represents the RGB image sequence captured over the interval $(t_{n}-t_{1})$ by the user-centric camera (Resolution: $1408\times 1408$, 10 FPS). $\mathcal{G}\in\mathbb{R}^{H_{2}\times W_{2}\times(t_{n}-t_{1})}$ corresponds to the grayscale eye-tracking image sequence (Resolution: $320\times 240$ per eye, 10 FPS), independently capturing the left and right eyes for gaze estimation.
The human-view inputs are utilized to infer the referred object, represented as $\overline{\mathcal{Z}_{g}}=(\overline{\mathbf{p}_{g}},\ \overline{\beta_{g}},\ \overline{M_{g}}),$ which encodes its spatial and visual properties. Here, $\overline{\mathbf{p}_{g}}\in\mathbb{R}^{2}$ denotes the 2D position of the referred object, $\overline{\beta_{g}}\in\mathbb{R}^{4}$ defines its bounding box, and $\overline{M_{g}}\in\mathbb{R}^{H_{1}\times W_{1}}$ represents its segmentation mask.
III-C Formulation of Robot View Representation
The robot camera captures the scene from the robot perspective, with input $\mathcal{C}\in\mathbb{R}^{H_{3}\times W_{3}}$ (Resolution: 1280 $\times$ 720) representing the image of the RGBD camera. The goal of the robot is to correlate the same referenced object within the robot view and the human view. This reference object in robot view is denoted as $\overline{\mathcal{Z}_{r}}=(\overline{\mathbf{p}_{r}},\ \overline{\beta_{r}},\ \overline{M_{r}})$. Similar to the human view, $\overline{\mathbf{p}_{r}}$, $\overline{\beta_{r}}$ and $\overline{M_{r}}$ encode the object’s spatial and visual properties within the robot’s perspective.
Figure 2: Human view intention fusion and multi-view intention alignment
III-D Formulation of Control System
The control system is responsible for generating and executing actions based on the intention of the human and robot views. The system is manipulated with a predefined set of parameterized action primitives $\mathcal{A}(\Theta)=\left\{a_{i}(\theta_{i})\mid i\in\mathbb{R}\right\}$ and follows a parameterized planning policy to ensure adaptive and context-aware manipulation. In $\mathcal{A}(\Theta)$ each primitive action $a$ is parameterized by 2D position, object category, translation distance, and rotation angle. These action primitives form low-level and high-level control commands, enabling the robot to execute pick-up, placing, pushing, moving, and rotating actions in a structured manner.
The execution strategy follows a parameterized planning policy $\pi$, which is denoted as:
$\pi:X\mapsto\ \mathcal{A}\times\Theta.$ (1)
As the input to the policy $\pi$, $X$ represents the state space, incorporating scene observations and task-specific constraints. $\mathcal{A}\times\Theta$ denotes a action-parameter pairs. $\Theta$ represents the parameters of action primitives. The policy generation will be discussed in Sec. IV-C.
We propose a framework for multi-modal HRI combined language and gaze via foundation models. Our proposed system diagram is shown in Fig. 1.
IV Methodology
Our system enables multi-modal HRI by integrating gaze, language, and visual perception across both human and robot viewpoints to achieve accurate intention inference and action execution. The methodology is structured into three key components: Human View Intention Fusion, Multi-View Intention Alignment, and Planning Policy Generation.
IV-A Human View Intention Fusion
When a command is given, the human’s gaze naturally shifts and fixates on the target only at specific moments within the command, making direct fixation-based selection unreliable. Our system aims to minimize this noise and achieve robust intention fusion by aligning gaze and language.
IV-A1 GPT-Agent for Human View Command Processing
To effectively fuse gaze and language for intention inference, we leverage a GPT-based agent (GPT-4o, frozen weight 2024-08-06, temperature = 1) to analyze the spoken command in text and determine the object or location referenced and the most relevant fixation period. This GPT-Agent takes $\mathcal{S}$ as input and produces the following structured output as:
$\mathbf{O}_{1}=\left\{\mathcal{L},\ c_{target},\ \Delta t\right\},$ (2)
where $\mathcal{L}$ defines the target properties, with possible values “object” or “position”. The variable $c_{target}$ specifies the target object category in the command, later utilized for scene observation in both human and robot views. If the referenced is a general category, the output is “stuff”; if it specifies a position without an object, it is labeled “position”; otherwise, it corresponds to a specific object category. $\Delta t$ indicates the time period of the word spoken by the user determined by Fast Whisper [29], during which the user is most likely fixating on the target object or location. For instance, given the command: “please put the apple there on the table”. The system interprets it as: $\mathbf{O}_{1}=\left\{(object,\ position),\ (apple,\ table),\ (\Delta t_{apple},\ \Delta t_{there})\right\}$. This structured output aligns gaze fixation with spoken language to enhance intention inference.
IV-A2 Scene Observation in Human View
We define a scene observation set in human view $\mathcal{Z}_{g}$ to describe target object categories spatial and visual properties as $\mathcal{Z}_{g}=\left\{c_{target},\ \{\mathbf{p}_{g_{i}},\ \beta_{g_{i}},\ M_{g_{i}}\}_{i=1}^{N}\right\}$. In this set $\mathbf{p}_{g_{i}}\in\mathbb{R}^{2}$, $\beta_{g_{i}}\in\mathbb{R}^{4}$ and $M_{g_{i}}\in\mathbb{R}^{H_{1}\times W_{1}}$ are the 2D position, bbox and mask of the target object $i$ and $c_{target}$ is the category for target object from the first GPT-Agent. $\mathbf{p}_{g_{i}}$ is defined as the midpoint of the $\beta_{g_{i}}$. A total of $N$ objects of this category are observed. In our proposed approach, Grounding Dino SAM2 [30, 31] performs object detection, classification, and segmentation with prompt $c_{target}$ from Eq. 2.
IV-A3 3D Gaze Estimation and Reconstruction
For 3D gaze estimation, the ARIA glasses incorporate two eye-tracking cameras. We employ the SocialEye model222github.com/facebookresearch/projectARIA_eyetracking to predict the 3D gaze vector ${}^{gp_{t_{i}}}\mathbf{P}_{{t_{i}}}\in\mathbb{R}^{3}$ at time $t_{i}$ in the pupil frame of the glasses. To reconstruct the 3D gaze vector within $\Delta t$ at time $t_{i+n}$, donated as ${}^{gc_{t_{i+n}}}\mathbf{P}_{{t_{i+n}}}$, onto the frame ()$\prescript{gc}{}{(\cdot)}$ at time $t_{i}$, denoted as ${}^{gc_{t_{i}}}\mathbf{P}_{{t_{i+n}}}\in\mathbb{R}^{3}$, and then project them onto image plane ${}^{{t_{i}}}\mathcal{U}$ at $t_{i}$ as gaze point ${}^{gc_{t_{i}}}\mathbf{p}^{gaze}_{{t_{i+n}}}\in\mathbb{R}^{2}$ we use the transformations:
${}^{gc_{t_{i+n}}}\mathbf{P}_{{t_{i+n}}}$ $={}^{gc}T_{gp}{}^{gp_{t_{i+n}}}\mathbf{P}_{{t_{i+n}}}$
${}^{gc_{t_{i}}}\mathbf{P}_{{t_{i+n}}}$ $=({}^{s}T_{gc_{t_{i}}})^{-1}({}^{s}T_{gc_{t_{i+n}}}){}^{gc_{t_{i+n}}}\mathbf{P}_{{t_{i+n}}}$ (3)
${}^{gc_{t_{i}}}\mathbf{p}^{gaze}_{{t_{i+n}}}$ $=\mathcal{K}{}^{gc_{t_{i}}}\mathbf{P}_{{t_{i+n}}}$
[…]
IV-A4 Human View Intention Alignment
After projecting gaze vectors onto a single image as $\mathbf{p}^{gaze}_{{t_{i+n}}}$ and generating the corresponding scene observation $\mathcal{Z}_{g}$, the next goal is to fuse gaze and language to infer the specific referred object $\overline{\mathcal{Z}_{g}}$ of the user. In our proposed approach, this intention fusion is determined by a weighted distance function by calculating the distance among each target object to all the gaze points, denoted as:
$\overline{\mathcal{Z}_{g}}=\mathop{\arg\min}\limits_{\mathbf{p}_{g_{i}}\in\mathcal{Z}_{g}}\sum_{n=0}^{N}e^{\alpha(n-N)}\left\|{}^{gc_{t_{i}}}\mathbf{p}^{gaze}_{{t_{i+n}}}-\mathbf{p}_{g_{i}}\right\|$ (4)
Where $N=\left\lfloor\Delta t\times 20-1\right\rfloor$ and the weighting factor $e^{\alpha(n-N)}$ ensures that more recent gaze points contribute more to the final intention fusion.
IV-B Multi-View Intention Alignment
Our system incorporates two cameras: one user-centric camera on the glasses and a robot camera. Through human view intention fusion, the referred object $\overline{\mathcal{Z}_{{g}}}$ is determined, while the robot also needs to determine the referred object $\overline{\mathcal{Z}_{{r}}}$. Instead of relying on complex extrinsic calibration between the robot camera and the user-centric camera on the glasses, we implemented a feature-based alignment approach. We first obtain the scene observation of the robot view with the same process used for the human view, generating the scene observation set $\mathcal{Z}_{r}=\left\{c_{target},\ \{\mathbf{p}_{r_{i}},\ \beta_{r_{i}},\ M_{r_{i}}\}_{i=1}^{N}\right\}$. To establish a correspondence between the human and robot viewpoints, we restrict feature matching from keypoints inside the bounding box of $\overline{\mathcal{Z}_{g}}$ to $\mathcal{C}$ with Superglue [33]. The highest density of the matched keypoints within each bounding box in $\mathcal{Z}_{r}$ determines the most likely corresponding object. The matched object is denoted by:
$\overline{\mathcal{Z}_{{r}}}=\mathop{\arg\max}\limits_{\beta_{r_{i}}\in\mathcal{Z}_{r}}\sum_{\mathbf{k}_{j}\in\overline{\beta_{g}}}\mathbbold{1}(\mathbf{k}_{j}\in\beta_{r_{i}})\vspace{-5pt}$ (5)
Where $\mathbbold{1}(\cdot)$ is an indicator function that equals one if the matched keypoint $\mathbf{k}_{j}$ falls within the segmentation mask $\beta_{r_{i}}$. This ensures robust object alignment across the two viewpoints. The general process of human intention fusion and multi-view intention alignment is illustrated in Fig. 2.
IV-C Planning Policy Generation
The final stage of our proposed system involves generating a planning policy to execute the human intention. In our proposed system, the state space $X$ in Eq. 1 is defined as $X=\left\{\overline{\mathcal{Z}_{r}},\ \mathbf{O}_{1},\ \mathcal{S}\right\}$ including referred object, auxiliary from the first GPT-Agent and user’s command.
To translate this state space into an executable sequence of parameterized action primitives, we employ a GPT-based planning policy generator. We adapt GPT-4o with frozen checkpoints from 2024-08-06, using a temperature of 1 and input $X$ for policy generation. The model is prompted to generate semantic-level parameterized action primitives in JSON format, denoted as $\pi$.
To generate reasonable policy, we design structured prompts that enforce strict constraints on parameterized action primitives usage, the coordinate system, and the expected output format. By providing task descriptions, parameterized action primitive definitions, example tasks, and structured planning steps, we ensure that the LLM outputs consistent, interpretable, and physically feasible actions, reducing ambiguity and enhancing execution reliability. An example input and output of the planning policy generation is shown as follows:
Input:
$\mathbf{O}_{1}$: $(object,\ position),(apple,\ table),(\Delta t_{apple},\ \Delta t_{there})$
$\overline{\mathcal{Z}_{r}}$: $(Position_{apple},\ Position_{there})$
$\mathcal{S}$: $please\ put\ the\ apple\ there\ on\ the\ table$
Output:
$\pi$: $[open\ gripper],\ [pick,\ [^{\prime}apple^{\prime},\ position_{apple}]]$
   $,\ [close\ gripper],\ [put,\ [^{\prime}table^{\prime},\ position_{there}]]$
   $[open\ gripper]$
V Experiment Setup
V-A Perception and Manipulator Setup
Figure 3: Meta ARIA Glasses: Sensor Setup & Reference Frames
Visual inputs for human view are collected through the Meta ARIA glasses [^3] https://www.projectARIA.com/, which also captures auditory signals through the built-in microphone as shown in Fig. 3. The robot view is collected through the Intel Realsense D435i RGBD camera. Our system has been tested with individuals in real-world environments, as illustrated in Fig. 4 (a-d). The scene includes a Franka Emika Panda 7DOF robot manipulator and various manipulation objects. An Intel D435i RGBD camera is positioned to monitor the workspace, ensuring clear visibility of both the environment and the robot. Multi-modal data processing is performed on a computer equipped with an RTX 4080 GPU, utilizing 6105 MB of GPU memory during interaction.
[…]
Figure 4: Overview of Experimental Scenarios. Scenario 1 evaluates object selection among similar items. Scenarios 2-4 assess real-world robotic task execution at increasing complexity.
V-B Description of Experimental Scenarios
To evaluate our system, we designed four experimental scenarios as shown in Fig. 4. Our system achieves an average LLM inference latency of ca. 1600 ms per command, with a worst-case latency of ca. 4200 ms. A full breakdown of inference latency under different conditions is provided in the supplementary materials.
  1. 1.
    Object Selection Among Similar Items: In the scenario 1 ($S_{1}$) shown in Fig. 4 (a), the test area contains nine visually similar pawns, each with a distance of 15cm between them, requiring the user to select a specific one.
  2. 2.
    Task Execution in a Complex Environment:
    As shown in Fig. 4 (b-d), Scenarios 2-4 ($S_{2}$-$S_{4}$) evaluate the execution of robotic tasks in increasingly complex real-world environments. The test area contains two similar cups, two similar exotic fruits (unknown name), and two similar plates. We categorize tasks into three levels of complexity: 1) Single-Step Action ($S_{2}$): Tasks requiring a single, direct action and where only one target object is referred to, such as picking up an object. 2) Multi-Step Action ($S_{3}$): Sequential actions requiring multiple steps to complete a task and where two target objects are referred to, such as put an object in a plate. 3) Causal Action ($S_{4}$): Tasks where an earlier action influences a later one, involving more than two target objects. For example, placing a fruit on a plate, then pouring liquid from a cup onto it without explicitly mentioning the plate’s position again.
Moreover, a user study was conducted to evaluate user satisfaction and experience. For the experiments and user study, we recruited 32 individuals, including 18 from Germany, and 14 from Singapore, all of whom are fluent in English.
V-C Evaluation Metrics
To quantitatively assess system performance, we use two key metrics: success rate and interaction time. The success rate is defined as the proportion of correctly executed tasks ($N_{\text{correct}}$) to the total number of trials ($N_{\text{total}}$), given by $(N_{\text{correct}}/N_{\text{total}})*100\%$. Interaction time represents the efficiency of the system, measured as the time from when the user begins issuing a command to when the system completes interaction. This metric eliminates variations caused by different robot platforms, network situation, and computing configurations, ensuring a fair comparison across methods. The instruction of our proposed FAM-HRI and the survey can be found at Github.
VI Results and Discussion
VI-A Baseline Selection
In our proposed FAM-HRI, we compare our proposed FAM-HRI with the various SOTA interaction methods utilizing different modalities. CaP [21] and ProgPrompt [22] are language-only interaction, making $S_{1}$ inefficient, as users would be required to specify the exact row and column of the pawns. GesSentence [4] utilizes gesture-only for interaction, HapGloves [34] employs a wearable tactile gloves for users to teleoperate the robot, while PhysicalDemo [35] leverages impedance-based feedback, where the user physically guides the robot through specific tasks. NVP-HRI [18] combines language and deictic posture for interaction, while GIRAF [19] fuses language with deictic gestures to achieve efficient object selection. These approaches are efficient in selecting similar objects, however, they rely on body movements and are less convenient for users with mobility or motor disabilities. FreeView [10] enables robot teleoperation through a virtual GUI controlled by the user’s gaze. Additionally since Gesture Sentence use pre-trained supervised networks for environment observation therefore cannot deal with unfamiliar objects in the scene. The results of baseline comparison, depicted in Tab. I.
Due to the lack of open-source code or similar hardware devices, we were unable to reproduce [10] and [34]. Instead, we directly report their success rate of task completion and interaction time from the original papers. However, since these methods used teleoperation, the comparison of success rate of task completion is not entirely fair and should be interpreted as a reference only.
TABLE I: Comparison of different HRI approaches. Best results are bolded to support our contribution claims.
Method Interaction Modality Hands Free? Voice Free? Similar Object? Unfamiliar Object? Success Rate $\uparrow$ (%) Interaction Time $\downarrow$ (s)
$S_{1}$ $S_{2}$ $S_{3}$ $S_{4}$ $S_{1}$ $S_{2}$ $S_{3}$ $S_{4}$
ProgPrompt [22] Language 12 73 52 16 4.7 2.4 6.4 9.7
CaP [21] Language 36 100 $96$ 84 4.7 3.7 6.9 11.5
GesSentence [4] Gesture 67 82 73 68 5.4 5.1 10.2 22.8
HapGloves [34] Haptic - 100 100 100 - - - -
PhysicalDemo [35] Impedance 100 100 61 52 6.3 5.8 12.3 31.7
FreeView [10] Gaze - 100 100 100 - 55 90 180
NVP-HRI [18] MultiModal 88 100 93 86 $2.3$ 2.2 4.6 11.2
GIRAF [19] MultiModal 94 100 $\underline{98}$ $\underline{91}$ $\underline{2.1}$ $\underline{1.6}$ $\underline{2.9}$ $\underline{6.8}$
FAM-HRI (Ours) MultiModal $\underline{96}$ 100 100 $\underline{94}$ 1.8 1.3 2.2 5.9
VI-B Object Selection Among Similar Items
In the $S_{1}$, as described in Sec. VB and illustrated in Fig. 4 (a), all 32 participants received verbal instructions on how to use FAM-HRI. To evaluate the our FAM-HRI distinguishing between multiple similar objects, each participant completed ten trials of pawn selection as instructed. The results, presented in Tab. I, show that our system achieved the second-highest success rate among all baseline methods in $S_{1}$, only lower than Physical Demo [35], which requires the user to physically guide the robotic arm near the target object. However, FAM-HRI significantly outperformed gesture-only and language-only methods, demonstrating its reliability in challenging and cluttered environments.
VI-C Task Execution in Complex Environment
In $S_{2}-S_{4}$, described in Sec. VB and illustrated in Fig. 4 (b-d), the same 32 participants were asked to complete three levels of tasks using objects in the scene ($S_{2}$: Single-Step Action, $S_{3}$: Multi-Step Action, $S_{4}$: Causal Action). Each participant performed ten trials per task level.
As shown in Tab. I, FAM-HRI outperformed all baseline methods (except for the two baseline FreeView and HapGloves using teleoperation), achieving the highest success rate across all task levels, demonstrating its robustness in complex real-world scenarios. Our success rate exceeds that of GIRAF [19] and NVP-HRI [18], both of which rely on an RGB-D camera to capture the 3D human skeleton, which is prone to sensor noise and ambient occlusion that reduces the overall success rate of the task execution.
VI-D Comparison of Effectiveness
The results in Tab. I present the average interaction time across all participants for each of the four scenarios. Our proposed FAM-HRI achieves the shortest interaction time among all baseline methods, demonstrating its efficiency in processing seamless multi-modal inputs. Unlike language-only methods, which require precise verbal descriptions, or gesture-based approaches, which depend on physical movement, FAM-HRI leverages gaze-language fusion to streamline intent inference. Furthermore, methods relying on teleoperation or physical demonstration exhibit longer interaction times due to manual corrections and additional user input steps. These results confirm that FAM-HRI enables faster interactions, making it a practical and efficient solution for real-world human-robot collaboration.
VI-E Qualitative Field Survey on User Experience
To rigorously assess the proposed method, we conducted a user survey to evaluate interaction experience, including participants with mild upper-limb impairments. Participants responded to questions using a five-point Likert scale, assessing FAM-HRI and baseline methods in terms of object selection accuracy, device flexibility, approach modernity, ease of use, and interaction efficiency. As shown in Fig. 5, users rated FAM-HRI higher than all baselines across qualitative measures. Notably, it outperformed baselines in efficiency and accuracy, with all participants favoring FAM-HRI for its seamless integration of gaze and language signals. Additionally, most users found our system more modern and easier to operate, emphasizing its intuitive design. The hands-free interaction enabled by ARIA glasses significantly reduced physical effort, offering a more accessible and user-friendly solution for human-robot collaboration, particularly benefiting individuals with limited mobility.
Figure 5: Results of User Experience Survey. The answer of the 32 participants are centered around zero for better comparison.
VII Conclusion and Future works
We proposed FAM-HRI, a multi-modal human-robot interaction framework that fuses real-time gaze and speech inputs from ARIA glasses with robot vision using GPT-based agents. Our system integrates human intention fusion, which identifies the target object and the relevant fixation period, with a multi-view alignment strategy that reconciles observations from the human and robot perspectives. The planning policy generation module then produces robust, parameterized action primitives for precise and efficient task execution in dynamic and cluttered environments. Experimental evaluations and user studies confirm the improved interaction performance offered by our approach.
However, FAM-HRI has several limitations. The computational latency introduced by LLMs inference and VLM may affect real-time responsiveness, making deployment on embedded systems or low-compute devices challenging. Additionally, speech input noise may introduce errors in object selection. A significant source of error arises from loose-fitting ARIA glasses, which can lead to misalignment in gaze estimation. Furthermore, our multi-view alignment relies on SuperGlue-based feature matching, but its robustness under object occlusion has not been extensively analyzed, potentially limiting performance in complex scenes.
Future work will focus on real-time optimization, exploring lighter models for edge deployment, and incorporating additional sensor modalities to enhance system robustness and adaptability in diverse HRI scenarios.

### v2

@@
 Effective Human-Robot Interaction (HRI) is crucial for enhancing accessibility and usability in real-world robotics applications.
-However, existing solutions often rely on gestures or language commands, making interaction inefficient and ambiguous, particularly for users with physical impairments.
+However, existing solutions often rely on gestures-only or language-only commands, making interaction inefficient and ambiguous, particularly for users with physical impairments.
 In this paper, we introduce FAM-HRI, an efficient multi-modal framework for human-robot interaction that integrates language and gaze inputs via foundation models.
@@
 Experimental evaluations demonstrate that FAM-HRI achieves a high success rate in task execution while maintaining a low interaction time, providing a practical solution for individuals with limited physical mobility or motor impairments.
+To support the community, we will release our system design, algorithms, and solutions as open-source after paper acceptance.
+Note to Practitioners
+In the field of assistive and service robotics, enabling users with physical disabilities to communicate intentions naturally remains a significant challenge.
+Existing uni-modal approaches, such as voice-only or gaze-only interaction, often struggle with ambiguity and environmental interference, limiting their reliability in real-world scenarios.
+Current multi-modal systems typically process gaze and speech separately, making it difficult to align user’s visual attention with spoken commands accurately.
+Additionally, many systems rely on bulky hardware or require users to maintain prolonged focus, reducing practicality and comfort.
+To address these issues, this paper proposes a foundation-model-assisted multi-modal human-robot interaction (FAM-HRI) framework that combines lightweight AR glasses with LLMs to fuse gaze and speech inputs in real time.
+The system automatically filters noisy gaze signals, aligns them with spoken commands, and generates precise robot actions.
+Experiments demonstrate that this approach achieved task success rates exceeding 94 % across multiple scenarios and outperformed all baseline methods in interaction efficiency, with the shortest interaction times in comparative evaluations.
+This method is well-suited for applications such as household assistance, and collaborative tasks where hands-free interaction is desirable.
+In the future, we plan to optimize computing performance and explore strategies to resolve conflicts between gaze and speech inputs, further enhancing usability and adaptability in diverse environments.
+Index Terms:
+Robot vision system, Robot control, Perception, Cooperative systems.
 {strip}
-[…]
-figureVoice-gaze fusion HRI enables efficient, intuitive interaction without memorization, ideal for physically impaired users.
-Supplementary Material
-The supplementary materials and video can be found at: https://github.com/laiyuzhi/FAM-HRI
+Figure 1: Voice-gaze fusion HRI enables efficient, intuitive interaction without memorization, ideal for physically impaired users.
 I Introduction
-Individuals with limited mobility or motor impairments, such as neuromuscular disorders, often struggle with daily tasks.
-These tasks can be automated using robotic manipulators, enhancing independence and improving quality of life [1, 2, 3].
-However, improving life through automation requires effective and accessible Human-Robot Interaction (HRI), which remains a challenge.
-Ensuring smooth interaction is particularly difficult in systems that rely on verbal commands [4].
-Accurately processing verbal descriptions becomes problematic when identifying unfamiliar objects or distinguishing between similar items.
-Existing voice-controlled assistants, such as Amazon Alexa or Tmall Genie, remain insufficient in bridging this gap for embodied HRI.
-Gaze, as a natural interaction modality, has great potential in HRI for individuals with severe physical limitations [5, 6, 7, 8], as shown in Fig.
-FAM-HRI: Foundation-Model Assisted Multi-Modal Human-Robot Interaction Combining Gaze and Speech.
-However, existing methods often rely on external or bulky devices, restricting mobility and reducing the accuracy of HRI [5, 7, 8].
-Some early gaze-based interfaces [6, 9] enable robot control at the cost of increased interaction time and complexity [10].
-Beyond hardware constraints, the nature of gaze detection itself adds challenges [11].
-During natural gaze interactions, the gaze does not always remain fixed on the target object but instead briefly lingers around the area [12].
-This dynamic nature increases noise in target selection process, reducing gaze stability in daily use.
-These challenges emphasize the need for accurate, efficient, and lightweight systems to enhance assistive robotics.
-To address the challenges, we introduce a Foundation-model Assisted Multi-modal HRI framework, FAM-HRI, that enables users to convey their intentions to robots.
-By wearing lightweight glasses from the Meta ARIA Research Kit [13], the system can simultaneously collect gaze and voice signals from users.
-By integrating these modalities, the robot accurately selects target objects and executes intended actions based on environmental context, reducing ambiguity in language-based systems, especially for users with limited mobility.
-A command-processing LLM analyzes user input to identify the optimal gaze period, extracting high-quality signals while minimizing noise from gaze dynamics.
-A motion-planning LLM then generates the robot’s action sequences, leveraging advanced reasoning to handle complex contexts more effectively than traditional rule-based systems.
+In recent years, assistive robotics has attracted increasing attention due to its potential to enhance the autonomy and quality of life of individuals with physical disabilities or neuromuscular impairments.
+Tasks that most people perform effortlessly—like reaching for a cup or opening a drawer—can be daunting for users with limited motor control.
+Robotic manipulators offer a promising path to automate these actions [1, 2, 3].
+However, realizing this vision hinges on effective and inclusive Human-Robot Interaction (HRI).
+Traditional interfaces, such as joysticks or touchscreen panels, demand dexterous control and are thus often inaccessible.
+Voice-controlled systems are more inclusive, yet remain insufficient in accurately interpreting users’ intentions, particularly when referring to unfamiliar objects or choosing among similar items [4, 5].
+Additionally popular voice-driven assistants like Amazon Alexa or Tmall Genie lack embodied grounding and often fail in real-world complex tasks.
+Thus, there is a pressing need for more intuitive, context-aware interaction paradigms that can bridge the gap between human intent and robot execution.
+Recent works [6, 7, 8, 9] have explored alternative interaction modalities such as eye gaze.
+Gaze is a natural, low-effort cue that can reveal intent without requiring physical movement.
+Early gaze-driven systems have demonstrated feasibility in robot control [7, 10], yet they rely on bulky external trackers or head-mounted cameras that impede mobility and resist integration into daily environments [6, 8, 9].
+Moreover, combining gaze with traditional Graphical User Interface (GUI) elements increases interaction time and complexity [11].
+While some approaches attempt to fuse gaze with speech, they typically process the two modalities independently, missing critical cross-modal correlations.
+As a result, most systems fail to deliver both the precision and ease-of-use needed for everyday assistive HRI, especially in cluttered and dynamic environments of real-world homes.
+Despite its potential, leveraging gaze and speech in HRI introduces several fundamental challenges: (1) Natural gaze behavior is dynamic and noisy: the user’s eyes do not remain fixed but instead shift subtly around the target area [12], introducing temporal noise and uncertainty in object selection.
+Gaze patterns may reflect intention only briefly, complicating efforts to extract meaningful cues.
+(2) Verbal commands are often ambiguous—especially when users must describe unknown or visually similar items without using precise labels [5].
+(3) Existing systems generally require expensive or immobile hardware setups, which hinder deployment in daily living environments [6, 8].
+(4) Few methods model the temporal alignment between gaze and speech.
+In real-world scenarios, a user may look at an object moments before or after issuing a verbal command.
+Without a principled way to synchronize and interpret these signals, errors propagate and degrade the robot’s performance.
+These challenges underscore the need for lightweight, intelligent, and modality-coordinated interaction frameworks.
+To overcome these limitations, we propose FAM-HRI—a Foundation-model-Assisted Multi-modal Human-Robot Interaction framework that enables intuitive communication for users with limited mobility.
+Our system leverages lightweight wearable glasses from the Meta ARIA Research Kit [13], which can simultaneously capture egocentric video, gaze signals, and voice commands.
+A foundation language model agent processes the spoken command and selects the most semantically aligned gaze period, filtering out irrelevant or noisy frames caused by microsaccades or transient fixations.
+Another large model agent performs high-level motion planning based on multi-modal context and environmental cues.
+As a result, users can simply glance at their intended object and speak naturally—the system then robustly identifies the target and commands the robot to act.
+This joint modeling of gaze and speech enables more precise, adaptive, and accessible HRI, particularly in complex scenes.
+By reducing ambiguity and minimizing user effort, FAM-HRI bridges the gap between intent and robotic autonomy in assistive settings.
 Our main contributions are summarized below:
-  • To the best of our knowledge, FAM-HRI is the one of the first frameworks to integrate voice and gaze with a lightweight wearable device for human-robot collaboration.
+  • To the best of our knowledge, FAM-HRI is one of the first frameworks to integrate simultaneous voice and gaze with a lightweight wearable device for efficient human-robot collaboration.
+Unlike prior works that treats gaze and speech as separate inputs [11, 14], our proposed unified multi-modal HRI framework improves robustness and precision by aligning gaze and speech commands, enabling reliable disambiguation in complex environments.
+  • Our system optimizes gaze fixation timing and robot planning by integrating language, gaze, and environmental context.
+This ensures precise user intent alignment with robot actions, enhancing task execution.
+The experimental results illustrate that users can complete complex tasks with a high success rate with only brief instruction.
+  • Compared with the state-of-the-art HRI methods, FAM-HRI achieves higher interaction success rates and faster input processing.
 By leveraging ARIA glasses, our approach facilitates natural and intuitive communication, enhancing accessibility and ease of use.
 We assess its effectiveness through a comprehensive user experience study.
-  • Our system optimizes gaze fixation timing and robot planning by integrating language, gaze, and environmental context.
-This ensures precise user intent alignment with robot actions, enhancing task execution.
-  • FAM-HRI was benchmarked against state-of-the-art HRI methods, achieving higher interaction success rates and faster input processing.
-To support the community, we will release our system design, algorithms, and solutions as open-source at github.com/laiyuzhi/FAM-HRI.
-II Related Work
+II Related Works
 With advancements in computer vision and highly integrated hardware, robots are increasingly embedded in daily life, assisting with tasks from home services [15, 16] to industrial automation [17, 18].
@@
 An optimal system should enable unrestricted gaze-based interaction by accurately estimating user attention [11].
+Recent work such as [14] focuses on robust gaze-based intention prediction in real-world scenarios.
+Their method applies OPTICS clustering to filter out redundant gaze points, enabling robust intention prediction.However, this approach presents several limitations.
+First, it is not integrated with an actual robot control system, and thus lacks validation in closed-loop interaction.
+Second, it does not consider the alignment between gaze and language input; users can only select target objects through gaze, limiting natural multi-modal expression.
 In [25], researchers introduced a gaze-following model that predicts gaze direction and visual attention using RGB images from a robot’s perspective combined with YOLO-based object detection.
@@
 This enhances the ability of the robot to interpret user intentions, improving efficiency and naturalness in human-robot collaboration, particularly in industrial settings.
+In addition, some researches in [27] proposed a human-in-the-loop multi-modal fusion method that dynamically adjusts the weights of gaze, gesture, and speech modalities based on user feedback.
+This adaptive mechanism improves the system’s robustness by resolving conflicting inputs during interaction.
+However, the approach is limited to a fixed set of predefined object categories and supports only simple, hard-coded robot actions.
+As a result, it lacks the flexibility to engage in natural, open-ended interaction and cannot support more complex or extended task sequences.
 However, these approaches also require users to fixate on targets continuously, which feels unnatural compared to typical communication, where the gaze naturally shifts.
@@
 Recently, researchers introduced LaMI [29] to explore how LLMs can improve multi-modal HRI by reasoning over diverse inputs.
-However, this approach primarily focuses on backend reasoning and overlooks frontend measurement uncertainties and other external factors affecting perception and decision-making in real-world interactions.
+While LaMI provides a strong backend reasoning capability, it primarily focuses on symbolic-level interpretation based on human pose estimation, dialogue history and scene detection.
+This approach overlooks frontend measurement uncertainties and other external factors affecting perception and decision-making in real-world interactions.
+For example, LaMI infers the user’s attention direction based on head orientation to determine whether the user is referring to the robot or the workspace, rather than determining the specific object to which the user is visually attending.
 Unlike previous HRI approaches that struggle with frontend uncertainties and real-world perception challenges, our method leverages lightweight ARIA glasses to capture high-precision gaze data and speech signals.
@@
 Additionally, a second LLM agent analyzes scene context and user intent to generate optimal robot actions, ensuring more efficient and natural HRI.
-III Problem Definition
 Our proposed system integrates environmental perception from both the human and the robot points of view to facilitate effective collaboration.
 The human view is mainly used for intention fusion, capturing user-centric cues, while the robot view focuses on manipulation, allowing precise interaction with objects.
-To structure multi-modal signals for foundation model reasoning, we need to define sets of geometric frames and control parameters.
+III Problem Definition
+Our proposed system integrates environmental perception from both human and robot views to facilitate effective collaboration.
+The human view is mainly used for intention fusion, capturing user-centric cues, while the robot view focuses on manipulation, allowing precise interaction with objects.
+To structure multi-modal signals for foundation model reasoning, we formalize sets of geometric frames and control parameters.
+For comprehensive definitions for the mathematical notations used throughout our paper, please refer to Appendix C.
 III-A Definitions of Frames and Transformations
@@
 The robot base frame ()$\prescript{r}{}{(\cdot)}$ serves as the primary reference for the robot, while the robot camera frame ()$\prescript{c}{}{(\cdot)}$ is used for manipulation.
-The ARIA glasses camera frame (()$\prescript{gc}{}{(\cdot)}$) provides a first-person user-centric perspective.
-The ARIA pupil frame (()$\prescript{gp}{}{(\cdot)}$), centered at the midpoint of the glasses’ bridge, serves as the reference for gaze vector estimation.
+The ARIA glasses camera frame ()$\prescript{gc}{}{(\cdot)}$ provides a first-person user-centric perspective.
+The ARIA pupil frame ()$\prescript{gp}{}{(\cdot)}$, centered at the midpoint of the glasses’ bridge, serves as the reference for gaze vector estimation.
 The transformation $T\in\mathbb{SE}(3)$ denotes a rigid body transformation in the special Euclidean group.
@@
 Here, $\mathcal{S}\in\Sigma^{*}$ denotes the transcribed speech sequence.
-The audio is recorded at a 48 kHz sampling rate using the built-in microphone of the ARIA glasses and converted into text via Fast Whisper [29], with each word timestamped.
+The audio is recorded at a 48 kHz sampling rate using the built-in microphone of the ARIA glasses and is converted into text via Fast Whisper [30], with each word timestamped.
 Six embedded microphones improve speech clarity by suppressing background noise and improving transcription accuracy even in noisy environments.
-Additional details are provided in the supplementary material.
 $\mathcal{U}\in\mathbb{R}^{H_{1}\times W_{1}\times(t_{n}-t_{1})}$ represents the RGB image sequence captured over the interval $(t_{n}-t_{1})$ by the user-centric camera (Resolution: $1408\times 1408$, 10 FPS).
@@
 Here, $\overline{\mathbf{p}_{g}}\in\mathbb{R}^{2}$ denotes the 2D position of the referred object, $\overline{\beta_{g}}\in\mathbb{R}^{4}$ defines its bounding box, and $\overline{M_{g}}\in\mathbb{R}^{H_{1}\times W_{1}}$ represents its segmentation mask.
+$\overline{\mathbf{p}_{g}}$ is defined as the midpoint of the $\overline{\beta_{g}}$.
 III-C Formulation of Robot View Representation
@@
 Similar to the human view, $\overline{\mathbf{p}_{r}}$, $\overline{\beta_{r}}$ and $\overline{M_{r}}$ encode the object’s spatial and visual properties within the robot’s perspective.
-Figure 2: Human view intention fusion and multi-view intention alignment
 III-D Formulation of Control System
@@
 The system is manipulated with a predefined set of parameterized action primitives $\mathcal{A}(\Theta)=\left\{a_{i}(\theta_{i})\mid i\in\mathbb{R}\right\}$ and follows a parameterized planning policy to ensure adaptive and context-aware manipulation.
-In $\mathcal{A}(\Theta)$ each primitive action $a$ is parameterized by 2D position, object category, translation distance, and rotation angle.
+In $\mathcal{A}(\Theta)$, each primitive action $a$ is parameterized by 2D position, object category, translation distance, and rotation angle.
 These action primitives form low-level and high-level control commands, enabling the robot to execute pick-up, placing, pushing, moving, and rotating actions in a structured manner.
@@
 IV-C.
-We propose a framework for multi-modal HRI combined language and gaze via foundation models. Our proposed system diagram is shown in Fig. 1.
 IV Methodology
@@
 The methodology is structured into three key components: Human View Intention Fusion, Multi-View Intention Alignment, and Planning Policy Generation.
+Our proposed system diagram is shown in Fig.
+2.
 IV-A Human View Intention Fusion
@@
 The variable $c_{target}$ specifies the target object category in the command, later utilized for scene observation in both human and robot views.
-If the referenced is a general category, the output is “stuff”; if it specifies a position without an object, it is labeled “position”; otherwise, it corresponds to a specific object category.
+If the referenced target is a general category, the output is “stuff”; if it specifies a position without an object, it is labeled “position”; otherwise, it corresponds to a specific object category.
 $\Delta t$ indicates the time period of the word spoken by the user determined by Fast Whisper [30], during which the user is most likely fixating on the target object or location.
@@
 IV-A2 Scene Observation in Human View
+Figure 3: Schematic of 3D gaze estimation and reconstruction.
 We define a scene observation set in human view $\mathcal{Z}_{g}$ to describe target object categories spatial and visual properties as $\mathcal{Z}_{g}=\left\{c_{target},\ \{\mathbf{p}_{g_{i}},\ \beta_{g_{i}},\ M_{g_{i}}\}_{i=1}^{N}\right\}$.
@@
 A total of $N$ objects of this category are observed.
-In our proposed approach, Grounding Dino SAM2 [30, 31] performs object detection, classification, and segmentation with prompt $c_{target}$ from Eq.
+In our proposed approach, Grounding DINO [31] and SAM 2 [32] perform object detection, classification, and segmentation with prompt $c_{target}$ from Eq.
 2.
@@
 We employ the SocialEye model222github.com/facebookresearch/projectARIA_eyetracking to predict the 3D gaze vector ${}^{gp_{t_{i}}}\mathbf{P}_{{t_{i}}}\in\mathbb{R}^{3}$ at time $t_{i}$ in the pupil frame of the glasses.
-To reconstruct the 3D gaze vector within $\Delta t$ at time $t_{i+n}$, donated as ${}^{gc_{t_{i+n}}}\mathbf{P}_{{t_{i+n}}}$, onto the frame ()$\prescript{gc}{}{(\cdot)}$ at time $t_{i}$, denoted as ${}^{gc_{t_{i}}}\mathbf{P}_{{t_{i+n}}}\in\mathbb{R}^{3}$, and then project them onto image plane ${}^{{t_{i}}}\mathcal{U}$ at $t_{i}$ as gaze point ${}^{gc_{t_{i}}}\mathbf{p}^{gaze}_{{t_{i+n}}}\in\mathbb{R}^{2}$ we use the transformations:
+As shown in Fig.
+3, to reconstruct the 3D gaze vector within $\Delta t$ at time $t_{i+n}$, donated as ${}^{gc_{t_{i+n}}}\mathbf{P}_{{t_{i+n}}}$, onto the frame ()$\prescript{gc}{}{(\cdot)}$ at time $t_{i}$, denoted as ${}^{gc_{t_{i}}}\mathbf{P}_{{t_{i+n}}}\in\mathbb{R}^{3}$, and then project them onto image plane ${}^{{t_{i}}}\mathcal{U}$ at $t_{i}$ as gaze point ${}^{gc_{t_{i}}}\mathbf{p}^{gaze}_{{t_{i+n}}}\in\mathbb{R}^{2}$ we use the transformations:
 ${}^{gc_{t_{i+n}}}\mathbf{P}_{{t_{i+n}}}$ $={}^{gc}T_{gp}{}^{gp_{t_{i+n}}}\mathbf{P}_{{t_{i+n}}}$
@@
 After projecting gaze vectors onto a single image as $\mathbf{p}^{gaze}_{{t_{i+n}}}$ and generating the corresponding scene observation $\mathcal{Z}_{g}$, the next goal is to fuse gaze and language to infer the specific referred object $\overline{\mathcal{Z}_{g}}$ of the user.
-In our proposed approach, this intention fusion is determined by a weighted distance function by calculating the distance among each target object to all the gaze points, denoted as:
+In our proposed approach, this intention fusion is determined by a weighted distance function by calculating the distance between each target object to all the gaze points, denoted as:
 $\overline{\mathcal{Z}_{g}}=\mathop{\arg\min}\limits_{\mathbf{p}_{g_{i}}\in\mathcal{Z}_{g}}\sum_{n=0}^{N}e^{\alpha(n-N)}\left\|{}^{gc_{t_{i}}}\mathbf{p}^{gaze}_{{t_{i+n}}}-\mathbf{p}_{g_{i}}\right\|$ (4)
-Where $N=\left\lfloor\Delta t\times 20-1\right\rfloor$ and the weighting factor $e^{\alpha(n-N)}$ ensures that more recent gaze points contribute more to the final intention fusion.
+Where $N=\left\lfloor\Delta t\times 20-1\right\rfloor$ and the weight factor $e^{\alpha(n-N)}$ ensures that more recent gaze points contribute more to the final intention fusion.
+Figure 4: Normalized Weights under Different $N$.
+Figure 5: Comparison of Intention Alignment with and without the Weight Factor.
+[…]
+  1. 1.
+    Ensuring Recent Gaze Points Have Higher Contribution: The exponential function $e^{\alpha(n-N)}$ weights recent gaze points higher while decaying older ones.
+  2. 2.
+    Avoiding Over-Reliance on a Single Gaze Point: The upper threshold of 0.65 prevents a single gaze point from dominating the selection.
 IV-B Multi-View Intention Alignment
@@
 The general process of human intention fusion and multi-view intention alignment is illustrated in Fig.
-2.
+6.
+Figure 6: Human view intention fusion and multi-view intention alignment.
+The user’s speech and gaze are aligned using a GPT-Agent, and the user intention is then processed by a multi-view alignment module to get the referred object within the robot manipulator view.
 IV-C Planning Policy Generation
@@
 The model is prompted to generate semantic-level parameterized action primitives in JSON format, denoted as $\pi$.
-To generate reasonable policy, we design structured prompts that enforce strict constraints on parameterized action primitives usage, the coordinate system, and the expected output format.
+In particular, our action primitives include both composite actions and atomic actions to flexibly support a wide range of task compositions:
+  1. 1.
+    Composite actions involve task-oriented behaviors commonly encountered in daily HRI scenarios:
+    1. (a)
+      Pick: grasp an object and lift it from the surface.
+    2. (b)
+      Put: place an object at a designated location.
+    3. (c)
+      Pour: tilt and empty the contents of a cup into a target container.
+    4. (d)
+      Swap: exchange the positions of two objects.
+    5. (e)
+      MoveTo: move the end-effector to a position.
+  2. 2.
+    Atomic actions offer fine-grained control over the robot’s motion and end-effector state:
+    1. (a)
+      MoveX, MoveY, MoveZ: move the end-effector along the x, y, or z axes.
+    2. (b)
+      OpenGripper: open the gripper.
+    3. (c)
+      CloseGripper: close the gripper.
+    4. (d)
+      Rotate: rotate the end-effector.
+To generate a reasonable policy, we design structured prompts that enforce strict constraints on parameterized action primitives usage, the coordinate system, and the expected output format.
 By providing task descriptions, parameterized action primitive definitions, example tasks, and structured planning steps, we ensure that the LLM outputs consistent, interpretable, and physically feasible actions, reducing ambiguity and enhancing execution reliability.
@@
 Input:
-$\mathbf{O}_{1}$: $(object,\ position),(apple,\ table),(\Delta t_{apple},\ \Delta t_{there})$
-$\overline{\mathcal{Z}_{r}}$: $(Position_{apple},\ Position_{there})$
-$\mathcal{S}$: $please\ put\ the\ apple\ there\ on\ the\ table$
+$\mathbf{O}_{1}$: $(object,\ position),(apple,\ table),(\Delta t_{apple},\ \Delta t_{here})$
+$\overline{\mathcal{Z}_{r}}$: $(Position_{apple},\ Position_{here})$
+$\mathcal{S}$: $please\ put\ the\ apple\ here\ on\ the\ table$
 Output:
@@
 Figure 7: Meta ARIA Glasses: Sensor Setup & Reference Frames
-Visual inputs for human view are collected through the Meta ARIA glasses [^3] https://www.projectARIA.com/, which also captures auditory signals through the built-in microphone as shown in Fig.
-3.
+Visual inputs for human view are collected through the Meta ARIA glasses [^3] https://www.projectARIA.com/, which also capture auditory signals through the built-in microphone.
+Examples of the outputs of the sensors used in the experiment and the reference frames are shown in Fig.
+7.
 The robot view is collected through the Intel Realsense D435i RGBD camera.
@@
 An Intel D435i RGBD camera is positioned to monitor the workspace, ensuring clear visibility of both the environment and the robot.
-Multi-modal data processing is performed on a computer equipped with an RTX 4080 GPU, utilizing 6105 MB of GPU memory during interaction.
-[…]
+Multi-modal data processing is performed on a computer equipped with a RTX 4080 GPU, utilizing 6105 MB of GPU memory during interaction.
 Figure 8: Overview of Experimental Scenarios. Scenario 1 evaluates object selection among similar items. Scenarios 2-4 assess real-world robotic task execution at increasing complexity.
 V-B Description of Experimental Scenarios
-To evaluate our system, we designed four experimental scenarios as shown in Fig.
-4.
-Our system achieves an average LLM inference latency of ca.
-1600 ms per command, with a worst-case latency of ca.
-4200 ms.
-A full breakdown of inference latency under different conditions is provided in the supplementary materials.
+To evaluate our system, we designed four experimental scenarios as shown in Fig. 8.
   1. 1.
     Object Selection Among Similar Items: In the scenario 1 ($S_{1}$) shown in Fig.
-4 (a), the test area contains nine visually similar pawns, each with a distance of 15cm between them, requiring the user to select a specific one.
+8 (a), the test area contains nine visually similar pawns, each with a distance of 15 cm between them, requiring the user to select one of the pawns designated by the experiment organizer by gazing at the bottom center of the pawns.
+The gaze error was reported in this scenario.
   2. 2.
@@
 We categorize tasks into three levels of complexity: 1) Single-Step Action ($S_{2}$): Tasks requiring a single, direct action and where only one target object is referred to, such as picking up an object.
-2) Multi-Step Action ($S_{3}$): Sequential actions requiring multiple steps to complete a task and where two target objects are referred to, such as put an object in a plate.
+2) Multi-Step Action ($S_{3}$): Sequential actions requiring multiple steps to complete a task and where two target objects are referred to, such as putting an object on a plate.
 3) Causal Action ($S_{4}$): Tasks where an earlier action influences a later one, involving more than two target objects.
@@
 Moreover, a user study was conducted to evaluate user satisfaction and experience.
-For the experiments and user study, we recruited 32 individuals, including 18 from Germany, and 14 from Singapore, all of whom are fluent in English.
+Fig.
+9 shows the experimental setup.
+For the experiments and user study, a total of 32 participants were recruited, comprising 18 participants from Germany and 14 from Singapore.
+All participants are affiliated with local universities and fluent in English.
+Among the German participants, none have physical impairments; the group includes 4 women and 14 men, with one individual over the age of 60 and the rest under 50.
+Among the Singaporean participants, 6 have mild upper-limb impairments; the group consists of 6 women and 8 men, with two individuals over the age of 60 and the rest under 50.
+Each individual received only brief instructions and was asked to naturally move their eyes among the multiple referred objects during the experiment.
+Furthermore, we formulated hypotheses regarding user per- ceptions (H1, H2), and performed hypothesis testing to evaluate statistical significance.
+Figure 9: Experimental Setup for user study. A disabled individual equipped with Aria glasses issues commands to the robot manipulator.
 V-C Evaluation Metrics
@@
 Interaction time represents the efficiency of the system, measured as the time from when the user begins issuing a command to when the system completes interaction.
-This metric eliminates variations caused by different robot platforms, network situation, and computing configurations, ensuring a fair comparison across methods.
-The instruction of our proposed FAM-HRI and the survey can be found at Github.
+This metric eliminates variations caused by different robot platforms, network situations, and computing configurations, ensuring a fair comparison across methods.
+In addition, we define task complexity as the sum of (1) the number of predefined actions in the action primitives and (2) the number of parameters, where parameters encompass both the number of referred objects and the associated spatial specifications such as positions and angles.
+We have also included a supplementary video that presents experimental results, illustrating the system’s performance across various tasks and environments.
 VI Results and Discussion
 VI-A Baseline Selection
-In our proposed FAM-HRI, we compare our proposed FAM-HRI with the various SOTA interaction methods utilizing different modalities.
-CaP [21] and ProgPrompt [22] are language-only interaction, making $S_{1}$ inefficient, as users would be required to specify the exact row and column of the pawns.
-GesSentence [4] utilizes gesture-only for interaction, HapGloves [34] employs a wearable tactile gloves for users to teleoperate the robot, while PhysicalDemo [35] leverages impedance-based feedback, where the user physically guides the robot through specific tasks.
+In our proposed FAM-HRI, we compare our FAM-HRI with the various SOTA interaction methods utilizing different modalities.
+CaP [21] and ProgPrompt [22] are language-only interactions, making $S_{1}$ inefficient, as users would be required to specify the exact row and column of the pawns.
+GesSentence [4] utilizes gesture-only interaction, HapGloves [35] employs wearable tactile gloves for users to teleoperate the robot, while PhysicalDemo [36] leverages impedance-based feedback, where the user physically guides the robot through specific tasks.
 NVP-HRI [5] combines language and deictic posture for interaction, while GIRAF [19] fuses language with deictic gestures to achieve efficient object selection.
-These approaches are efficient in selecting similar objects, however, they rely on body movements and are less convenient for users with mobility or motor disabilities.
+These approaches are efficient in selecting similar objects.
+However, they rely on body movements and are less convenient for users with mobility or motor disabilities.
 FreeView [11] enables robot teleoperation through a virtual GUI controlled by the user’s gaze.
-Additionally since Gesture Sentence use pre-trained supervised networks for environment observation therefore cannot deal with unfamiliar objects in the scene.
-The results of baseline comparison, depicted in Tab.
+Additionally, since Gesture Sentence use pre-trained supervised networks for environment observation, therefore cannot deal with unfamiliar objects in the scene.
+Comparison results are shown in Tab.
 I.
-Due to the lack of open-source code or similar hardware devices, we were unable to reproduce [10] and [34].
-Instead, we directly report their success rate of task completion and interaction time from the original papers.
-However, since these methods used teleoperation, the comparison of success rate of task completion is not entirely fair and should be interpreted as a reference only.
 TABLE I: Comparison of different HRI approaches. Best results are bolded to support our contribution claims.
@@
 GesSentence [4] Gesture 67 82 73 68 5.4 5.1 10.2 22.8
-HapGloves [34] Haptic - 100 100 100 - - - -
+HapGloves [35] Haptic 100 100 100 100 - - - -
 PhysicalDemo [36] Impedance 100 100 61 52 6.3 5.8 12.3 31.7
@@
 FAM-HRI (Ours) MultiModal $\underline{96}$ 100 100 $\underline{94}$ 1.8 1.3 2.2 5.9
-VI-B Object Selection Among Similar Items
+Remark: Due to the lack of open-source code for [11]. We directly report their success rate of task completion and interaction time from the original papers.
+“-” indicates that no relevant data are given explicitly or implicitly in the paper.
+VI-B Object Selection Among Similar Items and Gaze Error
 In the $S_{1}$, as described in Sec.
-VB and illustrated in Fig.
+V-B and illustrated in Fig.
 8 (a), all 32 participants received verbal instructions on how to use FAM-HRI.
-To evaluate the our FAM-HRI distinguishing between multiple similar objects, each participant completed ten trials of pawn selection as instructed.
+To evaluate our FAM-HRI distinguishing between multiple similar objects, each participant completed ten trials of pawn selection as instructed.
 The results, presented in Tab.
@@
 However, FAM-HRI significantly outperformed gesture-only and language-only methods, demonstrating its reliability in challenging and cluttered environments.
+We obtained the gaze error by measuring the distance between the gaze point and the center of the bottom of the pawn.
+With each participant wearing ARIA glasses correctly, the gaze error is $1.58\pm 0.62$ cm.
 VI-C Task Execution in Complex Environment
 In $S_{2}-S_{4}$, described in Sec.
-VB and illustrated in Fig.
+V-B and illustrated in Fig.
 8 (b-d), the same 32 participants were asked to complete three levels of tasks using objects in the scene ($S_{2}$: Single-Step Action, $S_{3}$: Multi-Step Action, $S_{4}$: Causal Action).
@@
 These results confirm that FAM-HRI enables faster interactions, making it a practical and efficient solution for real-world human-robot collaboration.
-VI-E Qualitative Field Survey on User Experience
+VI-E Evaluation of Tabletop Task
+TABLE II: Detailed simulation tabletop manipulation success rate (%) and interaction time (s) across different task scenarios
+Success Rate (%) Interaction time (s) Number of parameters Number of actions Complexity
+With referred object
+pick up the $<$object$>$ 100 1.3 1 1 2
+grab the pieces ($S_{1}$) 96 1.8 1 1 2
+put the $<$object$>$ on the $<$plate$>$ 100 2.2 2 3 5
+put the $<$object$>$ on the $<$plate$>$ then
+pour some thing from the $<$cup$>$ on it 94 5.9 4 6 10
+put this $<$object$>$ $<$there$>$ (Position on table) 89 2.5 2 3 5
+put the $<$object1$>$ and $<$object2$>$ on the $<$plate$>$ 92 3.8 3 6 9
+put the $<$object1$>$ on the $<$plate1$>$ then
+put the $<$object2$>$ on the $<$plate2$>$ 90 6.1 4 6 10
+grab the $<$object$>$ and lift up for $<$distance$>$ then
+turn it for $<$angle$>$ degrees 96 8.3 3 3 6
+Without referred object
+pick up this 100 1.1 1 1 2
+grab this ($S_{1}$) 87 0.8 1 1 2
+put this on that 98 1.7 2 3 5
+put this on that then pour something from this on it 96 4.2 4 6 10
+put this $<$there$>$ (Position on table) 93 1.9 2 3 5
+put this and this on that 96 3.1 3 6 9
+put this on this then put this on that 92 4.2 4 6 10
+grab this and lift it up for $<$distance$>$ then
+turn it for $<$angle$>$ degrees 95 7.9 3 3 6
+Remark: The listed commands are indicative examples.
+Our system supports a wide range of natural language expressions with similar semantic intention, allowing flexible and intuitive user interaction beyond fixed templates.
+Table II presents the success rates and average interaction times for various tabletop manipulation tasks.
+A key observation is that tasks executed without a referred object generally exhibit faster interaction times due to shorter command lengths.
+However, in Scenario $S_{1}$, the success rate for commands without explicit object references is lower.
+This is primarily because small, detailed parts of the chess pawns were sometimes misdetected by VLM.
+Consequently, during multi-view alignment, insufficient keypoints in these small detected parts were found, reducing the accuracy of object selection.
+Additionally, as the length and complexity of the user’s command increase, the reasoning process required by the two GPT-Agents becomes more involved, which in turn leads to a decrease in success rate due to higher inference uncertainty.
+VI-F Qualitative Field Survey on User Experience
 To rigorously assess the proposed method, we conducted a user survey to evaluate interaction experience, including participants with mild upper-limb impairments.
+All participants were asked to complete a survey at the end of the experiment.
 Participants responded to questions using a five-point Likert scale, assessing FAM-HRI and baseline methods in terms of object selection accuracy, device flexibility, approach modernity, ease of use, and interaction efficiency.
 As shown in Fig.
-5, users rated FAM-HRI higher than all baselines across qualitative measures.
+10, users rated FAM-HRI higher than all baselines across qualitative measures.
 Notably, it outperformed baselines in efficiency and accuracy, with all participants favoring FAM-HRI for its seamless integration of gaze and language signals.
 Additionally, most users found our system more modern and easier to operate, emphasizing its intuitive design.
-The hands-free interaction enabled by ARIA glasses significantly reduced physical effort, offering a more accessible and user-friendly solution for human-robot collaboration, particularly benefiting individuals with limited mobility.
-Figure 5: Results of User Experience Survey. The answer of the 32 participants are centered around zero for better comparison.
-VII Conclusion and Future works
+This user study investigated two main hypotheses:
+  1. 1.
+    H1: FAM-HRI is preferred by participants compared to all other baseline methods.
+  2. 2.
+    H2: FAM-HRI is quantitatively more effective and accurate compared to all other baseline methods.
+H1: Participants rated FAM-HRI higher than all baseline methods as “light and flexible” ($p<0.001$), “modern” ($p<0.001$), and “simple to use” ($p<0.001$).
+H2: Participants further rated the execution of their personal preferences, FAM-HRI, as “accurate” ($p<0.001$) and “efficient” ($p<0.001$).
+The results show that our proposed approach significantly reduced physical effort, offering a more accessible and user-friendly solution for human-robot collaboration.
+Figure 10: Results of User Experience Survey.
+The answers of the 32 participants are centered around zero for better comparison.
+Users rated FAM-HRI higher than all baselines across qualitative measures.
+VI-G Task Execution under varying noise levels
+Figure 11: System robustness evaluation under varying background noise. The background noise was introduced using a laptop. The system consistently achieved 100% success rate up to 80 dB.
+To evaluate the robustness of FAM-HRI in real-world conditions, we conducted experiments under varying ambient noise levels.
+As shown in Fig.
+11(a), the background noise was introduced using a laptop placed near the user, playing a pre-recorded dialogue video.
+The content of the video simulated a real indoor conversation scenario, ensuring that the noise characteristics are similar to those encountered in household or public settings.
+A sound level meter was positioned next to the user to measure the real-time noise intensity.
+The noise intensity ranged from 20 dB to 90 dB, covering common noise conditions, such as conversation, traffic, vacuum cleaner, and hairdryer.
+According to the Level Comparison Chart  [^4] http://ehs.yale.edu/sites/default/files/files/decibel-level-chart.pdf, hearing loss may result from sustained exposure to 90dB or more.
+The results, presented in Fig.
+11(b), illustrate that our proposed FAM-HRI maintains a high success rate even under increasing noise levels.
+Here, the user needs to complete a pick-placing task.
+The system consistently achieved 100% success rate up to 70 dB, indicating strong resilience to moderate noise conditions.
+Even at 90+ dB, where typical speech recognition systems often fail, FAM-HRI still maintained an 85% success rate, demonstrating its robustness in handling speech input under challenging conditions.
+The strong noise resistance is attributed to two aspects:
+  1. 1.
+    Hardware: Unlike traditional microphones that are directly exposed to ambient noise, the ARIA glasses’ microphones are embedded within the frame.
+This placement significantly reduces the impact of background noise, allowing the user’s voice to be clearly captured, even in high-noise environments.
+  2. 2.
+    Software: Our two LLM-Agent frameworks further enhance robustness against background noise.
+The GPT-Agent for human view command processing incrementally analyzes task-relevant verbs, nouns, and pronouns in the user’s command, filtering out irrelevant words and recognized background noise before passing a cleaned and structured instruction to the agent for policy generation.
+VI-H Latency of LLMs Inference
+Tab.
+III presents the inference latency of the two GPT-Agents under different experimental scenarios.
+The reported values include the average response time for human view command processing and planning policy generation, measured across all user trials.
+The results show that command processing latency varies between 1353ms and 2202ms, with increased latency in more complex scenarios due to additional language reasoning.
+Similarly, the policy generation latency ranges from 1281ms to 1997ms, reflecting the increasing complexity of task planning as the number of required action primitives grows.
+Despite these variations, our proposed FAM-HRI maintains a reasonable inference latency, demonstrating its feasibility for human-robot collaboration.
+TABLE III: Inference latency under different conditions
+Sceranio Latency of Human View Command Processing Latency of Policy Generation
+$S_{1}$ $1353\pm 753$ms $1281\pm 692$ms
+$S_{2}$ $1624\pm 660$ms $1535\pm 946$ms
+$S_{3}$ $1855\pm 725$ms $1803\pm 891$ms
+$S_{4}$ $2202\pm 936$ms $1997\pm 861$ms
+Remark: The reported latency values correspond to inference using a full-scale LLM without any quantization [37], distillation [38], or retrieval-augmented generation (RAG) [39].
+Incorporating such optimization strategies is expected to yield substantial reductions in latency.
+VII Conclusion
 We proposed FAM-HRI, a multi-modal human-robot interaction framework that fuses real-time gaze and speech inputs from ARIA glasses with robot vision using GPT-based agents.
@@
 Experimental evaluations and user studies confirm the improved interaction performance offered by our approach.
-However, FAM-HRI has several limitations.
+VIII Limitations and Future Works
+FAM-HRI has several limitations.
 The computational latency introduced by LLMs inference and VLM may affect real-time responsiveness, making deployment on embedded systems or low-compute devices challenging.
-Additionally, speech input noise may introduce errors in object selection.
-A significant source of error arises from loose-fitting ARIA glasses, which can lead to misalignment in gaze estimation.
-Furthermore, our multi-view alignment relies on SuperGlue-based feature matching, but its robustness under object occlusion has not been extensively analyzed, potentially limiting performance in complex scenes.
-Future work will focus on real-time optimization, exploring lighter models for edge deployment, and incorporating additional sensor modalities to enhance system robustness and adaptability in diverse HRI scenarios.
+Future work will explore trade-offs between accuracy and speed, such as fine-tuning parameter-efficient large models using Low-Rank Adaptation (LoRA) [40] and implementing retrieval-augmented generation (RAG) [41] with smaller backbones to reduce latency.
+A further significant source of error arises from loose-fitting ARIA glasses, which can lead to misalignment in gaze estimation.
+Cluttered or outdoor environments, where lighting variations, occlusions, and background motion introduce additional uncertainty.
+To address these limitations, future work will focus on developing a VLM that explicitly incorporates diverse environmental conditions during training [41], enabling the system to generalize more robustly across a wider range of real-world HRI scenarios.
+Moreover, our current framework does not explicitly handle rare but possible situations in which a user’s speech command and gaze trajectory refer to different objects—e.g., the user looks at a cup on the left while verbally referring to a similar cup on the right.
+In our current method, the robot naturally resolves such ambiguities by selecting the object with the lowest weighted distance to the gaze trajectory.
+While these conflicting cases are unlikely in practical settings, we acknowledge their existence and plan to address them in future iterations.
+One direction involves integrating a dialogue module [42] that allows the robot to perform clarification actions, such as hovering above the candidate object to verify user intention before task execution.
+This would enable the system to robustly disambiguate multi-modal commands in more complex or ambiguous contexts.
+Appendix A Failure Modes
+Figure 12: Failure Modes.
+As shown in Fig. 12, the failure cases of our system can be categorized into four key aspects: human view input, scene observation, multi-view alignment, and LLM reasoning.
+  1. 1.
+    Human View Input: For human view input, there are two sources of failure, speech recognition errors and inaccurate gaze estimation.
+The speech recognition error in our system primarily due to misinterpretation of similar-sounding words.
+Gaze estimation errors are largely influenced by improper wearing of the glasses.
+Specifically, when the nose bridge of the ARIA glasses does not correctly placed on the user’s nose, the accuracy of gaze estimation decrease significantly.
+However, to mitigate this, users were instructed to adjust their glasses properly before proceeding with the experiment, maximise ensuring that such errors do not propagate into the system.
+  2. 2.
+    Scene Observation: For scene observation, failures mainly come from more or less detected objects.
+In Scenario $S_{1}$, where users need to select a specific chess piece without explicitly stating its category, smaller parts of the chess pieces were sometimes incorrectly detected as separate objects.
+Additionally, speech recognition errors could result in the misclassification of the target object, causing the system to detect an incorrect category and subsequently select the wrong object.
+  3. 3.
+    Multi-View Alignment: For multi-view alignment, the main reason for failure is insufficient or unclear feature correspondence between human and robot views.
+Feature matching using superglue becomes unreliable in the presence of weak object textures, repetitive patterns, or partial occlusions, leading to incorrect correspondences.
+  4. 4.
+    LLM reasoning: For LLM Reasoning, the primary failure modes come from formatting inconsistencies and hallucinations in the generated output.
+Since FAM-HRI requires the LLM’s response to strictly follow a predefined prompt format, any deviation renders the output unusable by subsequent system modules.
+In rare cases, the LLM misinterprets the required action parameters, leading to incorrect parameter values or an incorrect ordering of action arguments.
+Additionally, logic errors happen in policy generation due to LLM hallucinations.
+For instance, in placement tasks, the generated policy may not open the gripper after reaching the target position, leading to task failure.
+However, such errors were observed infrequently in our experiments, as structured prompt design and output constraints significantly reduced the occurrence of these issues.
+Appendix B Parameter Tuning and Reproducibility
+TABLE IV: List of Parameters
+Grounding Dino SAM2
+Box Threshold 0.3
+Text Threshold 0.3
+Grounding-Model Tiny
+SAM2 Model Large
+SuperGlue
+Max Keypoints 10000
+Keypoints Threshold 1e-5
+Match Threshold 1e-5
+Resize No
+Weights indoor
+Our approach involves several parameters that are manually set or empirically tuned based on experimental observations.The Weight Factor in intention alignment was determined through extensive testing to balance recent gaze points’ influence while preventing a single point from dominating the selection process.
+While these parameters have been carefully adjusted to optimize system performance, we acknowledge that they are not mathematically proven to be optimal.
+To ensure reproducibility, we provide explicit parameter settings for key components of our system in Tab. IV, including Grounding Dino for Scene Observation and SuperGlue for Multi-View Alignment.
+Appendix C List of symbols and their meanings
+To provide a comprehensive reference for the mathematical notations used throughout our paper, we include Tab.
+V in the supplementary materials.
+This table details all symbols, spaces, sets, and corresponding meanings, covering essential aspects of frame transformations, gaze reconstruction, scene observations, intention fusion, control system variables, and policy generation.
+The notation system is structured to facilitate clarity in understanding multi-view geometric transformations, gaze-language fusion, and LLM-based policy generation. Specifically, it includes:
+  1. 1.
+    Frame representations for robot base, cameras, and gaze.
+  2. 2.
+    Scene Observations for both human and robot view, ensuring a unified representation for multi-view alignment.
+  3. 3.
+    Control system parameters, including action primitives, parameters, and planning policies.
+This structured notation serves as a foundation for understanding FAM-HRI’s multi-modal interaction system, enabling reproducibility and further development within the research community.
+TABLE V: List of symbols and their meanings
+Symbol Space Meaning
+()$\prescript{r}{}{()}$ Frame: Robot Base
+()$\prescript{c}{}{()}$ Frame: Robot Camera
+()$\prescript{gc}{}{()}$ Frame: Glasses Camera
+()$\prescript{gp}{}{()}$ Frame: Glasses Pupil
+$\mathcal{K}$ $\mathbb{R}^{3\times 3}$ Basic Geometry: Camera Intrinsic Matrix
+$\mathbf{p}$ $\mathbb{R}^{2}$ Basic Geometry: 2D Position of a 3D coordinate
+$\mathbf{P}$ $\mathbb{R}^{3}$ Basic Geometry: 3D coordinate
+${}^{gc}T_{gp}$ $\mathbb{SE}(3)$ Transformation from Glasses Camera Frame to Glasses Pupil Frame
+${}^{s}T^{head}_{t_{i}}$ $\mathbb{SE}(3)$ Transformation Representing Head Pose at Time $t_{i}$
+$\beta$ $\mathbb{R}^{4}$ Environmental Bounding Box Representation $\left([x_{\min},y_{\min},x_{\max},y_{\max}]\right)$
+$M$ $\mathbb{R}^{H\times W}$ Environmental Segmentation Mask (binary or probability map)
+$\mathcal{S}$ $\Sigma^{*}$ Human Input Speech Sequence in Text
+$\mathcal{G}$ $\mathbb{R}^{H\times W}$ Human input of Eye tracking camera
+$\mathcal{U}$ $\mathbb{R}^{H\times W*(t_{n}-t_{1})}$ Human view: Image from User-Centric Camera on glasses
+$\mathcal{Z}_{H}$ $\Sigma^{*}\times\mathbb{R}^{H_{1}\times W_{1}\times(t_{n}-t_{1})}\times\mathbb{R}^{H_{2}\times W_{2}\times(t_{n}-t_{1})}$ Human Input State Vector / Descriptor
+${}^{gc_{t_{i}}}\mathbf{P}_{{t_{i}}}$ $\mathbb{R}^{3}$ Human view: 3D Gaze Vector at Time $t_{i}$ in Glasses Carmea Frame at Time $t_{i}$
+${}^{gc_{t_{i}}}\mathbf{P}_{{t_{i+n}}}$ $\mathbb{R}^{3}$ Human view: 3D Gaze Vector at Time $t_{i+n}$ in Glasses Carmea Frame at Time $t_{i}$
+$\mathcal{Z}_{g}$ Human view: Scene Observations Set $\left\{c_{target},\ \{\mathbf{p}_{g_{i}},\ \beta_{g_{i}},\ M_{g_{i}}\}_{i=1}^{N}\right\}$
+$\mathbf{p}^{gaze}$ $\mathbb{R}^{2}$ Human View: 2D Projected Gaze Point on the Image Plane of $\mathcal{U}$
+${}^{gc_{t_{i}}}\mathbf{p}^{gaze}_{{t_{i+n}}}$ $\mathbb{R}^{2}$ Human View: Projected Gaze Point at Time $t_{i+n}$ on Image at Time $t_{i}$
+$\overline{\mathcal{Z}_{g}}$ $(\overline{\mathbf{p}_{g}},\ \overline{\beta_{g}},\ \overline{M_{g}})$ Human view: Referred Object
+$\mathcal{C}$ $\mathbb{R}^{H\times W}$ Robot view: Image from Camera of Robot
+$\mathcal{Z}_{r}$ Robot view: Scene Observations Set $\left\{c_{target},\ \{\mathbf{p}_{r_{i}},\ \beta_{r_{i}},\ M_{r_{i}}\}_{i=1}^{N}\right\}$
+$\overline{\mathcal{Z}_{r}}$ $(\overline{\mathbf{p}_{r}},\ \overline{\beta_{r}},\ \overline{M_{r}})$ Robot view: Referred Object
+$\mathbf{O}_{1}$$\left\{\mathcal{L},\ c_{target},\ \Delta t\right\}$ Control system: LLM Output for Gaze Time Period Prediction
+$\mathcal{L}$ $\Sigma^{*}$ Control System: Target Property Descriptor (e.g., Object, Position)
+$c_{target}$ $\Sigma^{*}$ Control System: Target Object Category from User Command
+[…]
+$X$$\left\{\overline{\mathcal{Z}_{r}},\ \mathbf{O}_{1},\ \mathcal{S}\right\}$ Control System: State Space
+$\mathbf{O}_{2}$ $\pi$ Control System: LLM Output for Policy Generation
+$\theta$ $\mathbb{R}^{n}$ Control System: Action Parameter Space (Rotation, Translation, position, etc.)
+$a$ $\Sigma^{*}$ Control System: Action Space Primitives
+$\mathcal{A}$$\left\{a(\theta)\right\}$ Control System: Set of Action Primitives
+[…]
+$\pi$ $\pi:X\mapsto\ \mathcal{A}\times\Theta$ Control System: Parameterized Planning Policy for Executing Actions